# Dialogue Without Limits: Constant-Sized KV Caches for Extended Responses in LLMs

### v1

@@ Abstract
Autoregressive Transformers rely on Key-Value (KV) caching to accelerate inference. However, the linear growth of the KV cache with context length leads to excessive memory consumption and bandwidth constraints. This bottleneck is particularly problematic in real-time applications – such as chatbots and interactive assistants – where low latency and high memory efficiency are critical. Existing methods drop distant tokens or compress states in a lossy manner, sacrificing accuracy by discarding vital context or introducing bias.
We propose MorphKV, an inference-time technique that maintains a constant-sized KV cache while preserving accuracy. MorphKV balances long-range dependencies and local coherence during text generation. It eliminates early-token bias while retaining high-fidelity context by adaptively ranking tokens through correlation-aware selection. Unlike heuristic retention or lossy compression, MorphKV iteratively refines the KV cache via lightweight updates guided by attention patterns of recent tokens. This approach captures inter-token correlation with greater accuracy, crucial for tasks like content creation and code generation. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works, enabling efficient real-world deployment.
@@ 1 Introduction
Numerous approaches have been proposed in the literature to minimize the impact of growing KV cache sizes. In studies like FlashAttention (Dao, 2023) and vLLM (Kwon et al., 2023), the authors propose techniques to either materialize only partial caches at a time or use paging techniques by fragmenting KV caches into smaller blocks, thereby avoiding the need to reserve memory for the entire cache at once. Beyond this, prior works like FastGen (Ge et al., 2023), and MiKV (Yang et al., 2024) compress KV caches by retaining only a subset of KV pairs from recent and older tokens, prioritizing those deemed important based on attention scores and discarding the rest. However, this creates a trade-off: while memory savings increase as more KVs are discarded, the accuracy depends on the retained KVs effectively capturing context for future tokens. Consequently, these methods often sacrifice accuracy for reduced memory usage.
@@
Keyformer (Adnan et al., 2024) selects top old and recent tokens for memory savings and uses Gumbel noise to reduce selection bias. While somewhat effective, it cannot fully eliminate selection bias because Gumbel noise by itself introduces new forms of biases (Mussmann et al., 2017), reducing accuracy for long-context and long-response tasks. SnapKV (Li et al., 2024), the current state-of-the-art, achieves high accuracy in long-context tasks by retaining the most attended tokens from the input prompt. However, as shown in Figure 2(a), it retains all generated output tokens from the decode phase, causing the KV cache size to scale with response lengths, making it unsuitable for long-response tasks. Overcoming these limitations is crucial for improving LLM performance in applications like scripting and content creation (long-response tasks).
@@
To better capture context, MorphKV prioritizes the attention scores of relevant recent tokens rather than relying on historically most-attended tokens, addressing bias issues observed in methods like H2O and Keyformer. As shown in Figure 2(b), MorphKV achieves better scores than SnapKV: for Phi4, up to 8% higher for long-context task (VCSum) while saving 56% on KV cache memory, and for Qwen2.5, up to 21% higher score for long-response task (LongGenBench) while saving 83% on KV cache memory. This shows the impact of retaining a compact set of high-quality KVs and an improved attention mechanism proposed in MorphKV.
@@ 2.2 Limitations of KV Cache Compression Methods
KV cache compression addresses their growing memory footprint through quantization, algorithmic optimizations, cross-layer approaches, and pruning. Quantization-based methods store KV pairs using lower precision (Kang et al., 2024; Zhang et al., 2024), whereas algorithmic methods modify attention-layer computations (Chang et al., 2024; Saxena et al., 2024). Cross-layer optimizations leverage inter-layer similarities, selectively retaining KVs from layers with significant contributions (Yuan et al., 2024; Saxena et al., 2024; Cai et al., 2024). Pruning strategies selectively retain a subset of important KV pairs (Adnan et al., 2024; Xiao et al., 2023; Li et al., 2024; Zhang et al., 2023; Liu et al., 2024b) and are more effective than other approaches because they capture task-specific context more accurately.
However, compressed KV caches are either limited by accuracy or scalability. Constant-sized KV caching methods, such as Scissorhands, StreamingLLM, and Keyformer, have limited accuracy. On the other hand, more accurate methods like SnapKV are not scalable as they fail to address growing KV cache size for long-response tasks.
@@ 3.3 Handling Multiple Heads
For transformer-based LLMs, attention is performed across multiple heads, also referred to as Multi-Headed Attention (MHA)  (Vaswani et al., 2023). In MHA with $M$ heads, each head maintains a separate KV cache, resulting in $M$ distinct caches and significant memory overhead. Modern LLMs like Llama3.1  (Dubey et al., 2024) use Grouped-Query Attention (GQA) (Ainslie et al., 2023) with $M^{\prime}$ grouped-key heads ($M^{\prime}<M$), where every $M/M^{\prime}$ query heads share a single key/value head, reducing the number of KV caches to $M^{\prime}$ and cutting memory usage by a factor of $M/M^{\prime}$.
@@ 5.1.1 Performance
Performance is measured using an LLM-based Judge (Mistral-Large-123B), which assigns scores based on several criteria, and these scores are aggregated to form a final overall score. Table 1 shows that MorphKV outperforms both H2O, and SnapKV on Llama, Mistral and Phi4 models, while achieving comparable performance to SnapKV for Qwen. Notably, MorphKV consistently scores higher than H2O on metrics such as relevance (see Appendix A.2), indicating its ability to track important context for long responses when constrained by memory size.
@@
Table 1: LongWriter: LLM Judge Score comparison of MorphKV, H2O, SnapKV, and Full-Attention scores across different models. MorphKV outperforms other techniques by upto $4.5\%$, retaining important older tokens at a much smaller memory footprint.
@@ 5.1.2 KV cache Sizes
Figure 6 shows the normalized KV cache sizes for H2O, SnapKV, and MorphKV relative to Full-Attention. On average, MorphKV needs only 0.25×0.25$\times$0.25 × the memory capacity of Full-Attention, while H2O needs roughly 1$\times$ and SnapKV needs as much as 4$\times$ the memory capacity of Full-Attention.
Figure 6: LongWriter: KV cache sizes for H2O, SnapKV, MorphKV relative to Full-Attention. On average,  MorphKV requires 0.25$\times$ the KV cache size used by Full-Attention, while H2O and SnapKV use 1$\times$ and 4$\times$ respectively.
@@ 5.2 Long-Response: LongGenBench Tasks
LongGenBench contains structured long-response tasks for temporal and spatial categories. The temporal category is divided into Diary Entry and Menu planning tasks, while the spatial category includes Skyscraper Design and Urban Planning. The original dataset has 400 samples (100 from each sub-category). For a fair comparison under limited resources, we select 40 samples, with ten from each sub-category, and use greedy decoding capped at 8K tokens. SnapKV employs all attention heads with a 32-token window and a total KV cache capacity of 1K tokens per head, whereas H2O keeps 4K tokens in the cache, also maintaining all attention heads. In contrast, MorphKV uses a 200-token recent window and a 4K-token total capacity across GQA heads, adopting max() fusion due to its consistently higher performance than sum(). The larger window enables MorphKV to retain critical distant tokens during generation (refer to Section 6.2 for more information).
Table 2: LongGenBench: Performance comparison of MorphKV, SnapKV, H2O across different models. MorphKV achieves better scores across almost all evaluation metrics i.e., Completion Rate (CR), Accuracy Once, Accuracy Range, Accuracy Periodic, and Accuracy Average.
@@ 5.2.1 Performance
Table 2 shows the performance of MorphKV over prior works. LongGenBench uses a rigorous evaluation suite to assess response quality. This includes information recall about singular instances (Accuracy Once), range of instances (Accuracy Range), periodic instances (Accuracy Periodic) and their average (Average Accuracy), while Completion Rate (CR) quantifies the percentage of tasks successfully completed. MorphKV generally outperforms or matches SnapKV, and H2O on all models and metrics. Notably, SnapKV retains all prompt tokens due to its ample cache budget. It also keeps track of every decoded token, effectively replicating Full-Attention for these tasks.
@@ 5.2.2 KV cache Sizes
Figure 8 shows the KV cache sizes for H2O, SnapKV, and MorphKV relative to Full-Attention. On average, MorphKV requires 0.55×0.$5$ while H2O and SnapKV require 1.22×1.$2$, and 5.01×5.01$\times$5.01 × the cache size of Full-Attention respectively.
Figure 8: LongGenBench: KV cache usage of H2O, SnapKV, and MorphKV relative to Full-Attention. SnapKV has an explosive KV cache size (up to 13$\times$ higher) due to the extensive number of KV pairs retained. In contrast, MorphKV maintains a constant footprint, as it retains a fixed number of recent and older tokens.
@@
Figure 9: LongBench: (a) Llama3.1-8b-instruct KV cache sizes of SnapKV, and MorphKV relative to full attention. On an average, SnapKV has a KV cache size of 0.42×0.42$\times$0.42 ×, whereas MorphKV is 0.15×0.15$\times$0.15 × compared to Full-Attention (b) Average KV cache sizes of SnapKV and MorphKV relative to full attention across different models. MorphKV yields comparable performance to SnapKV at roughly 50% lower KV cache budget in a long-context setting, where the prompt is significantly larger than the response.
@@ 5.3 Long-Context: LongBench Tasks
Apart from being memory-efficient for long-response tasks, MorphKV also offers competitive performance as the state-of-the-art prompt KV compression for long-context tasks. We evaluate MorphKV, SnapKV, and Full-Attention across benchmarks in LongBench. LongBench is a comprehensive, bilingual, multitask benchmark suite used to evaluate LLMs for processing extended contexts. It comprises datasets across six task categories in English and Chinese, with an average prompt length of nearly 6K tokens. For MorphKV, we set the recent window to 32 tokens and fix its total cache capacity at 2K tokens, and use the sum() fusion. In contrast, SnapKV preserves 1024 tokens from the prefill phase and all decoded tokens across all attention heads.

### v2

@@ Abstract
Autoregressive Transformers rely on Key-Value (KV) caching to accelerate inference. However, the linear growth of the KV cache with context length leads to excessive memory consumption and bandwidth constraints. Existing methods drop distant tokens or compress states in a lossy manner, sacrificing accuracy by discarding vital context or introducing bias.
We propose MorphKV, an inference-time technique that maintains a constant-sized KV cache while preserving accuracy. MorphKV balances long-range dependencies and local coherence during text generation. It eliminates early-token bias while retaining high-fidelity context by adaptively ranking tokens through correlation-aware selection. Unlike heuristic retention or lossy compression, MorphKV iteratively refines the KV cache via lightweight updates guided by attention patterns of recent tokens. This approach captures inter-token correlation with greater accuracy, which is crucial for tasks like content creation and code generation. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works, enabling efficient deployment.
@@ 1 Introduction
Numerous approaches have been proposed in the literature to minimize the impact of growing KV cache sizes. In studies like FlashAttention (Dao, 2023) and vLLM (Kwon et al., 2023), the authors propose techniques to either materialize only partial caches at a time or use paging techniques by fragmenting KV caches into smaller blocks, thereby avoiding the need to reserve memory for the entire cache at once. Beyond this, prior works like FastGen (Ge et al., 2023) and MiKV (Yang et al., 2024) compress KV caches by retaining only a subset of KV pairs from recent and older tokens, prioritizing those deemed important based on attention scores and discarding the rest. However, this creates a trade-off: while memory savings increase as more KVs are discarded, the accuracy depends on the retained KVs effectively capturing context for future tokens. Consequently, these methods often sacrifice accuracy for reduced memory usage.
@@
Keyformer (Adnan et al., 2024) selects top old and recent tokens for memory savings and uses Gumbel noise to reduce selection bias. While somewhat effective, it cannot entirely eliminate selection bias because Gumbel noise by itself introduces new forms of biases (Mussmann et al., 2017), reducing accuracy for long-context and long-response tasks. SnapKV (Li et al., 2024), the current state-of-the-art, achieves high accuracy in long-context tasks by retaining the most attended tokens from the input prompt. However, as shown in Figure 2(a), it retains all generated output tokens from the decode phase, causing the KV cache size to scale with response lengths, making it unsuitable for long-response tasks. Overcoming these limitations is crucial for improving LLM performance in applications like scripting and content creation (long-response tasks).
@@
To better capture context, MorphKV prioritizes the attention scores of relevant recent tokens rather than relying on historically most-attended tokens, addressing bias issues observed in methods like H2O and Keyformer. As shown in Figure 2(b), MorphKV achieves better scores than SnapKV: for Phi4, up to 8% higher for long-context task (VCSum) while saving 56% on KV cache memory, and for Qwen2.5, up to 21% higher score for the long-response task (LongGenBench) while saving 83% on KV cache memory. This shows the impact of retaining a compact set of high-quality KVs and an improved attention mechanism in MorphKV.
@@
MorphKV is now open-source, and accessible at https://github.com/ghadiaravi13/MorphKV.
@@ 2.2 Limitations of KV Cache Compression Methods
KV cache compression addresses their growing memory footprint through several strategies, such as quantization, algorithmic optimizations, cross-layer and cross-head approaches, and pruning. Quantization-based methods store KV pairs using lower precision (Kang et al., 2024; Zhang et al., 2024), whereas algorithmic methods modify attention-layer computations (Chang et al., 2024; Saxena et al., 2024). Cross-layer optimizations leverage inter-layer similarities, selectively retaining KVs from layers with significant contributions (Yuan et al., 2024; Saxena et al., 2024; Cai et al., 2024). Cross-head optimization (Fu et al., 2024; Feng et al., 2024) reduces the KV cache footprint by retaining KVs only from the most impactful attention heads, as different heads contribute unevenly to model performance. Pruning strategies selectively retain a subset of important KV pairs (Adnan et al., 2024; Xiao et al., 2023; Li et al., 2024; Zhang et al., 2023; Liu et al., 2024b) and are more effective than other approaches because they capture task-specific context more accurately.
However, compressed KV caches are either limited by accuracy or scalability. Constant-sized KV caching methods, such as Scissorhands, StreamingLLM, and Keyformer, have limited accuracy. On the other hand, more accurate methods like SnapKV are not scalable as they fail to address the growing KV cache size for long-response tasks.
@@ 3.3 Handling Multiple Heads
For transformer-based LLMs, attention is performed across multiple heads, called Multi-Headed Attention (MHA)  (Vaswani et al., 2023). In MHA with $M$ heads, each head maintains a separate KV cache, resulting in $M$ distinct caches and significant memory overhead. Modern LLMs like Llama3.1  (Dubey et al., 2024) use Grouped-Query Attention (GQA) (Ainslie et al., 2023) with $M^{\prime}$ grouped-key heads ($M^{\prime}<M$), where every $M/M^{\prime}$ query heads share a single key/value head, reducing the number of KV caches to $M^{\prime}$ and cutting memory usage by a factor of $M/M^{\prime}$.
@@ 5.1.1 Performance
Performance is assessed using an LLM-based Judge (Mistral-Large-123B), with assigned scores across several criteria aggregated into a final metric. Table 1 shows that MorphKV outperforms H2O, and SnapKV on Llama, Mistral, and Phi4, while achieving comparable performance for Qwen. MorphKV consistently excels over H2O in relevance (see Appendix A.4), demonstrating its ability to retain critical context, even under memory constraints.
@@
Table 1: LongWriter: Comparison of LLM Judge Scores shows that MorphKV outperforms other methods by up to $4.5\%$, retaining important older tokens at a much smaller memory footprint.
@@ 5.1.2 KV cache Sizes
Figure 6 shows the normalized KV cache sizes relative to Full-Attention. On average, MorphKV reduces memory usage to 0.25×0.25$\times$0.25 × that of Full-Attention, while H2O requires 1$\times$ and SnapKV incurs significantly higher usage, up to 4$\times$.
Figure 6: LongWriter: KV cache usage relative to Full-Attention. On average,  MorphKV requires only 0.25$\times$ the KV cache size, while H2O and SnapKV consume 1$\times$ and 4$\times$ respectively.
@@ 5.2 Long-Response: LongGenBench Tasks
LongGenBench contains structured long-response tasks for temporal and spatial categories. The temporal category is divided into Diary Entry and Menu planning tasks, while the spatial category includes Skyscraper Design and Urban Planning. The original dataset has 400 samples (100 from each sub-category). For a fair comparison under limited resources, we select 40 samples, with ten from each sub-category, and use greedy decoding capped at 8K tokens. SnapKV employs all attention heads with a 32-token window and a total KV cache capacity of 1K tokens per head, whereas H2O keeps 4K tokens in the cache, also maintaining all attention heads. In contrast, MorphKV uses a 200-token recent window and a 4K-token total capacity across GQA heads, adopting max() fusion due to its consistently higher performance than sum(). The larger window enables MorphKV to retain critical distant tokens during generation (refer to Appendix A.2 for further information).
@@
Table 2: LongGenBench: Performance comparison of MorphKV, against SnapKV, H2O. MorphKV achieves better scores across all evaluation metrics i.e., Completion Rate (CR), and all grouped accuracy metrics (Accuracy Once, Range, Periodic, and Average).
@@ 5.2.1 Performance
Table 2 shows the performance of MorphKV over prior works. LongGenBench uses a rigorous evaluation suite to assess response quality. This includes information recall about singular instances (Accuracy Once), range of instances (Accuracy Range), periodic instances (Accuracy Periodic), and their average (Average Accuracy), while Completion Rate (CR) quantifies the percentage of tasks successfully completed. MorphKV generally outperforms or matches SnapKV, and H2O on all models and metrics. Notably, SnapKV retains all prompt tokens due to its ample cache budget. It also keeps track of every decoded token, effectively replicating Full-Attention for these tasks.
@@ 5.2.2 KV cache Sizes
Figure 8 shows the KV cache sizes for H2O, SnapKV, and MorphKV relative to Full-Attention. On average, MorphKV achieves significant memory savings, requiring only 0.55×0.$5$ the memory usage with Full-Attention, while H2O and SnapKV require 1.22×1.$2$, and upto 5.01×5.01$\times$5.01 × the cache size of Full-Attention, respectively.
Figure 8: LongGenBench: KV cache usage relative to Full-Attention. SnapKV incurs up to 13$\times$ higher due to the extensive retention of KV pairs, while MorphKV maintains a constant cache.
@@
Figure 9: LongBench: (a) Llama3.1-8B Instruct KV cache sizes of SnapKV, and MorphKV relative to full attention. On an average, SnapKV has a KV cache size of 0.42×0.42$\times$0.42 ×, whereas MorphKV is 0.15×0.15$\times$0.15 × compared to Full-Attention (b) Average KV cache sizes of SnapKV and MorphKV relative to full attention across different models. MorphKV yields comparable performance to SnapKV at roughly 50% lower KV cache budget in a long-context setting, where the prompt is significantly larger than the response.
@@ 5.3 Long-Context: LongBench Tasks
Besides being memory-efficient for long-response tasks, MorphKV also offers competitive performance as the state-of-the-art prompt KV compression for long-context tasks. We evaluate MorphKV, SnapKV, and Full-Attention across benchmarks in LongBench. LongBench is a comprehensive, bilingual, multitask benchmark suite used to evaluate LLMs for processing extended contexts. It comprises datasets across six task categories in English and Chinese, with an average prompt length of nearly 6K tokens. For MorphKV, we set the recent window to 32 tokens and fix its total cache capacity at 2K tokens, and use the sum() fusion. In contrast, SnapKV preserves 1024 tokens from the prefill phase and all decoded tokens across all attention heads.
@@ 5.3.2 KV cache Sizes
6 Runtime and System-Level Trade-offs
MorphKV incurs runtime overhead during response-generation as it introduces additional computation steps that depend on the attention profile of the recent window tokens, necessitating additional memory accesses. To minimize this overhead, the implementation of MorphKV employs dedicated CUDA streams to prefetch the attention profile ahead of time. By default, MorphKV performs token-eviction at each generation step to precisely manage KV cache memory usage, crucial for long-context and long-response tasks.
Table 4 shows that the memory savings with MorphKV significantly outweigh the runtime overhead, improving the overall system throughput. Furthermore, this runtime cost can be amortized by evicting tokens less frequently (see Appendix A.5), albeit at higher memory usage. These trade-offs highlight that KV cache compression techniques must navigate several system-level considerations, and optimizing for a single metric alone is inadequate for practical adoption.
xMetricx SnapKV H2O MorphKV
Runtime (lower is better) 1$\times$ 1.62$\times$ 1.50$\times$
Memory (lower is better) 1$\times$ 0.26$\times$ 0.14$\times$
Accuracy (higher is better) 1$\times$ 0.94$\times$ 1.01$\times$
Throughput (higher is better) 1$\times$ 2.40$\times$ 4.68$\times$
Table 4: LongBench: Despite introducing runtime overhead, MorphKV achieves substantial memory savings, leading to higher overall system throughput, while maintaining response accuracy
7 Discussion
7.1 Variation with Design Parameters
The design of MorphKV consists of three key design parameters, namely, fusion function used to create the attention profile, recent window size, and total KV cache budget. These hyperparameters collectively influence the overall performance. MorphKV shows minimal sensitivity to variations in its hyperparameters, highlighting its robustness across various configurations. We refer the reader to Appendix A.1A.2, and A.3 for detailed discussions on the sensitivity of MorphKV’s performance to the choice of fusion function, window sizes, and KV cache budget respectively.
7.2 Outperforming Full Attention
KV cache compression methods, despite relying on a subset of context tokens, often outperform full attention in practice (Li et al., 2024; Cai et al., 2024). This stems from their ability to focus on the most relevant tokens, reducing the influence of less useful ones. This is especially beneficial for long-response scenarios, where noise due to irrelevant tokens accumulates over time, leading to a noticeable decline in output quality, or degeneration (Holtzman et al., 2019). We examine the N-gram repetition rate in responses from the Llama3.1-8B Instruct model to measure this degeneration in generated responses. A higher repetition rate indicates higher degradation in output quality. Table 5 shows that MorphKV has the lowest repetition rate as it continues to prune the KV cache during token generation. In contrast, SnapKV and full attention exhibit greater redundancy as they retain many more tokens in the KV cache.
xMetricx MorphKV SnapKV Full-Attention
Repetition Rate (N=10) 68% 89% 89%
Table 5: LongWriter: Degeneration in Llama3.1-8B responses measured using N-gram (N=10) repetition. MorphKV reduces repetition by dynamically evicting tokens, limiting contextual noise.
8 Future Work
In addition to token selection, prior work has shown that KV cache can further be optimized across attention layers (Cai et al., 2024), and attention heads (Fu et al., 2024; Feng et al., 2024). MorphKV is complementary to these works, and integrating them would yield further memory savings. While a detailed evaluation of integrating MorphKV across other optimization axes is deferred to future work, we present a preliminary analysis with layer-aware considerations in Appendix A.6. Furthermore, to minimize the runtime cost with MorphKV, we plan to integrate MorphKV within the Flash Attention kernel, improving inference efficiency while preserving accuracy and memory benefits.
9 Conclusion
The growing memory footprint of KV caches in LLMs poses a critical bottleneck for long-context and long-response tasks. In this paper, we propose MorphKV that addresses this challenge by introducing a dynamic, correlation-aware token selection mechanism that maintains a constant-sized KV cache while preserving contextual coherence. MorphKV leverages attention profiles of recent tokens to retain only the most relevant distant tokens. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works. Our experiments demonstrate that MorphKV scales efficiently with response length, degrading only 10% in performance even as outputs grow to 12K tokens, compared to 15–18% degradation for state-of-the-art prior works. Furthermore, MorphKV’s compatibility with GQA enables 4$\times$ greater memory efficiency than MHA-based approaches, making it practical for real-world deployment. These advances position MorphKV as a practical inference-time solution, balancing accuracy and memory usage without sacrificing the ability to capture long-range dependencies.
Acknowledgments
This research has been supported by computing support on the Vista GPU Cluster through the Center for Generative AI (CGAI) and the Texas Advanced Computing Center (TACC) at the University of Texas at Austin. We thank the generous support from the Cockrell School of Engineering and the AMD endowment at the University of Texas at Austin. This work is supported in part by the Cisco Research Award. Prashant J. Nair is supported by Intel Transformation Server Architecture (TSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number RGPIN-2019-05059] Grants. We acknowledge Won Joon Yun for editorial feedback and Sourish Wawdhane for the code structuring effort.
Impact Statement
This work aims to advance the field of Machine Learning by reducing the memory footprint of large language model (LLM) inference while maintaining accuracy. By lowering hardware constraints, our technique potentially broadens the applicability of LLMs to more users and domains, including creative writing, data-intensive research, and education. However, it also amplifies existing concerns: easier access to powerful LLMs may exacerbate issues such as misuse of misinformation, large-scale automation of personalized or sensitive communication, and environmental impact due to increased computational usage. While the proposed method does not introduce new ethical risks beyond those already associated with LLMs, we encourage developers and practitioners to deploy it responsibly, carefully considering the contexts in which LLMs are employed and the societal implications of further democratizing access.
Appendix A Appendix
A.1 Impact of Fusion Function: sum() Versus max()
A.1.1 On Long-Context Tasks